### v1

# SSFO: Self-Supervised Faithfulness Optimization for Retrieval-Augmented Generation

Abstract
Retrieval-Augmented Generation (RAG) systems require Large Language Models (LLMs) to generate responses that are faithful to the retrieved context. However, faithfulness hallucination remains a critical challenge, as existing methods often require costly supervision and post-training or significant inference burdens. To overcome these limitations, we introduce Self-Supervised Faithfulness Optimization (SSFO), the first self-supervised alignment approach for enhancing RAG faithfulness. SSFO constructs preference data pairs by contrasting the model’s outputs generated with and without the context. Leveraging Direct Preference Optimization (DPO), SSFO aligns model faithfulness without incurring labeling costs or additional inference burden. We theoretically and empirically demonstrate that SSFO leverages a benign form of likelihood displacement, transferring probability mass from parametric-based tokens to context-aligned tokens. Based on this insight, we propose a modified DPO loss function to encourage likelihood displacement. Comprehensive evaluations show that SSFO significantly outperforms existing methods, achieving state-of-the-art faithfulness on multiple context-based question-answering datasets. Notably, SSFO exhibits strong generalization, improving cross-lingual faithfulness and preserving general instruction-following capabilities. We release our code and model at the anonymous link: https://github.com/chkwy/SSFO
1 Introduction
With the widespread deployment of Retrieval Augmented Generation (RAG) RAG ; jokinen2024need , large language models (LLMs) openai2023gpt ; touvron2023llama are increasingly expected to generate responses that adhere closely to the provided context trust_align ; rag-hat ; ragtruth , especially in scenarios where their internal parametric knowledge is insufficient or outdated. As a result, faithfulness hallucination zhou2023context ; survey_taxonomy ; es2024ragas has emerged as a critical challenge for current LLMs.
While a growing body of literature has emerged to address faithfulness hallucination, existing methods often suffer from two main drawbacks: reliance on costly supervision and post-training or increased inference burden. Current approaches generally fall into two categories: post-training-based methods and decoding strategy-based methods. Post-training-based methods trust_align ; rag-hat ; context-dpo ; liu2025chatqa leverage supervised fine-tuning and direct preference optimization dpo to enhance faithfulness. However, these methods often necessitate expensive human or advanced AI supervision (e.g., GPT-4). Meticulously creating thousands to tens of thousands of training data incurs significant annotation costs. On the other hand, decoding strategy-based methods gema2024decore ; shi2024trusting alleviate faithfulness hallucinations through a plug-and-play approach that can be easily adapted to newly developed LLMs. However, they typically double the inference burden by requiring parallel processing with perturbed and natural inputs, increasing computational complexity.
To overcome these limitations, we propose Self-Supervised Faithfulness Optimization (SSFO), the first self-supervised LLM faithfulness alignment approach. SSFO offers: 1) A self-supervised faithfulness alignment framework with negligible post-training cost (hundreds of self-generated examples), which helps SSFO easily adapt to newly developed LLM. 2) No added inference burden, which is crucial for lightweight deployment on edge devices yu2024edge . To implement SSFO, we leverage the model being optimized itself to generate preference data pairs by contrasting its outputs generated when provided with context with those generated from the query alone (i.e., without context). DPO dpo training is then applied using preference data pairs to align the model toward enhanced faithfulness. Our results demonstrate that this simple yet effective method can outperform both post-training-based methods trust_align ; context-dpo ; liu2025chatqa and decoding strategy-based methods gema2024decore in faithfulness to the context.
To understand the underlying mechanism of SSFO, we show that the improvement can be attributed to the likelihood displacement phenomenon LD . We provide both theoretical analysis and empirical results demonstrating that likelihood displacement transfers probability mass from parametric-based tokens to context-aligned tokens. Building on this insight, we propose a modified DPO loss function tailored to enhance this beneficial displacement. Overall, our contributions can be concluded as follows:
  • We introduce SSFO, the first self-supervised alignment method for LLM faithfulness alignment. SSFO leverages self-generated data during training, requiring no human annotations, strong AI models, or ground-truth labels; the training signal derives entirely from contrasting the model’s own parametric knowledge (as unpreferred examples) against retrieved knowledge (as preferred examples). We show that faithfulness alignment can be achieved through self-supervision without external supervision.
  • We provide a theoretical and empirical analysis of the underlying alignment mechanism of SSFO and introduce a modified DPO loss function to enhance the alignment process. We find that likelihood displacement can be beneficial in the RAG setting and that encouraging this displacement further boosts faithfulness alignment.
  • We provide comprehensive evaluations across diverse LLMs and benchmarks. Results show that SSFO achieves state-of-the-art faithfulness and superior generalization, including robust cross-lingual contextual faithfulness. Moreover, since trained with only hundreds of self-generated examples, SSFO preserves LLM’s general instruction-following ability, avoiding catastrophic forgetting catastrophic .
Figure 1: (a): Existing post-training methods rely on human annotators or strong AI models to build SFT or preference datasets, resulting in heavy labeling costs and lengthy post-training processes. (b) SSFO leverages the model itself to generate its own preference data: Given query $x$, it generates a context-grounded response $y_{c}^{\prime}$ (with external knowledge) and a parametric-only response $y_{p}$ (query only). SSFO can significantly reduce faithfulness hallucinations without any external supervision and incurs only negligible post-training costs.
2 Preliminaries
Faithfulness hallucination refers to the LLM’s response incorporating incorrect or made-up information that is not faithful to the provided context zhou2023context ; survey_taxonomy ; es2024ragas . Let $\pi_{\theta}$ represent the model with parameters $\theta$ which generates a response $y$ when provided with a query $x$ with provided context $c$. The response $y$ can be conceptualized as being composed of two components: $y_{c}$ corresponds to content derived from the external context $c$, whereas $y_{p}$ represents content originating from the model’s parametric knowledge. Specifically, response $y$ can be represented as the combination of these two components, i.e., $y=y_{c}\oplus y_{p}$. To construct a faithful response, we aim to minimize discrepancies between response $y$ and ideal context-grounded response $y_{c}$.
Directed Preference Optimization (DPO) dpo : RLHF is computationally expensivecheng2023adversarial ; yuan2023rrhf and can suffer from instabilitiessong2023reward ; go2023aligning . DPO can bypass both explicit reward estimation and performing reinforcement learning to learn the policy using a single maximum likelihood objective. The DPO loss is defined as:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right],$ (1)
where $(x,y_{w},y_{l})$ represents a data sample from dataset $\mathcal{D}$ consisting of a prompt $x$, a preferred completion $y_{w}$, and a unpreferred completion $y_{l}$. $\pi_{\theta}$ is the policy model undergoing optimization, and reference model $\pi_{\text{ref}}$ is the original state of the model before optimization. The hyperparameter $\beta$ controls the difference between policy model $\pi_{\theta}$ and reference model $\pi_{\text{ref}}$.
Likelihood Displacement LD ; pal2024smaug ; tajwar2024preference : Likelihood displacement is a counterintuitive phenomenon observed during direct preference optimization, where the probabilities for the preferred response $\pi_{\theta}(y_{w}\mid x)$ and the unpreferred response $\pi_{\theta}(y_{l}\mid x)$ both decrease, while the margin between them widens. Since $y_{w}$ is typically the optimal response (e.g., human-written or from a superior model), this reduction is problematic. Recent work dposhift ; AlphaPO aims to alleviate this phenomenon.
3 Methodology
In this section, we describe the methodology of the proposed self-supervised Faithfulness Optimization (SSFO). SSFO leverages self-supervised data construction and preference alignment training to reduce faithfulness hallucination in language models. Our goal is to train models to prioritize faithfulness to the provided external context over their internal parametric knowledge. This prioritization is critical for robust Retrieval-Augmented Generation (RAG) systems.
3.1 Self-Supervised Preference Data Construction
Existing approaches trust_align ; rag-hat ; context-dpo leverage DPO to mitigate faithfulness hallucinations and rely on curated preference data, often from human annotators or powerful AI models like GPT-4. While effective, these approaches incur substantial data annotation costs and post-training overhead.
To address this challenge, we propose a self-supervised data construction method that avoids any external labeling or supervision, as shown in Fig.˜1 (b). Our key idea is to exploit the LLM’s own responses under different knowledge access conditions to construct the preference pairs. Specifically, we generate two types of outputs for preference optimization:
Construction of preferred response $y_{w}$: We provide the model $\pi_{\theta}$ with both the query $x$ and the external context knowledge $c$ (from MS MARCO MS_MARCO ) to construct preferred responses, i.e., $y_{w}\sim\pi_{\theta}(y\mid x,c)$. Given the known faithful hallucination of LLMs (i.e. blend internal and external information when generating responses) trust_align ; ragtruth ; faithbench , we denote the response as a composition of context-based response $y_{c}$ and parametric-knowledge-based response $y_{p}$, such that $y_{w}=y_{c}^{\prime}=y_{c}\oplus y_{p}$.
Construction of unpreferred response $y_{l}$: We provide the model with the query $x$ only, omitting the external context. The model then generates a response based solely on its parametric knowledge: $y_{p}\sim\pi_{\theta}(y\mid x)$. We denote this response as $y_{l}=y_{p}$, since it reflects the model’s internal knowledge, and is more susceptible to hallucinations due to the absence of grounding in retrieved information.
The preference data pairs $(y_{c}^{\prime},y_{p})$ thus establish the context-grounded response as the positive example of faithfulness, and the purely parametric response as the negative example.
3.2 Self-Supervised Faithfulness Optimization
We perform Direct Preference Optimization over generated preference dataset $(y_{c}^{\prime},y_{p})$ to achieve faithfulness alignment. Specifically, Given a language model $\pi_{\theta}$, we minimize the following loss:
$\mathcal{L}\left(\pi_{\theta};\pi_{\mathrm{ref}}\right)=-\mathbb{E}_{\left(x,c,y_{c}^{\prime},y_{p}\right)\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{c}^{\prime}\mid x,c\right)}{\pi_{\mathrm{ref}}\left(y_{c}^{\prime}\mid x,c\right)}-\beta\log\frac{\pi_{\theta}\left(y_{p}\mid x,c\right)}{\pi_{\mathrm{ref}}\left(y_{p}\mid x,c\right)}\right)\right].$ (2)
This objective encourages the model to increase the likelihood of context-grounded response $y_{c}^{\prime}$ while penalizing the ungrounded, parametric-only response $y_{p}$. Since $y_{p}$ also contributes to the composite response $y_{c}^{\prime}$, this optimization inherently suppresses hallucinated content in $y_{c}^{\prime}$ and amplifies the influence of the grounded component $y_{c}$ within $y_{c}^{\prime}$. Through this formulation, we steer the model to reduce faithfulness hallucination (i.e. align more closely with retrieved context) by maximizing the discrepancy between outputs generated with external knowledge and those generated without it, without requiring costly external supervision.
In practice, training on just a few hundred such pairs yields significant improvements in faithfulness, outperforming methods that rely on human or strong AI-generated training data (Section˜4.3).
3.3 Analyzing Likelihood Displacement in Self-Supervised Faithfulness Optimization
Empirical studies Section˜4.1 show that although $y_{c}^{\prime}$ is an imperfect golden answer generated by $\pi_{\text{ref}}$, training with Eq.˜2 leads to a policy model $\pi_{\theta}^{*}$ that can significantly outperform $\pi_{\text{ref}}$. We attribute this performance improvement to a benign form of likelihood displacement LD ; pal2024smaug ; tajwar2024preference . Specifically, we demonstrate that in the context-based question-answering setting (i.e. RAG setting), SSFO facilitates this benign displacement by reallocating probability mass from tokens associated with parametric knowledge to those grounded in external contextual information.
s Refer to caption
Figure 2: Left: Log-likelihood of preferred response $\pi_{\theta}(y_{c}^{\prime}|x,c)$ versus unpreferred responses $\pi_{\theta}(y_{p}|x,c)$ over the course of SSFO optimization. Right: We compare the base model and optimized model on MemoTrap memotrap dataset and show the mean change in probability for context-based tokens $\Delta P(z_{c})$, and for parametric-only tokens $\Delta P(z_{p})$, revealing that optimization increases $\Delta P(z_{c})$ while decreasing $\Delta P(z_{p})$, $r$ denotes the Pearson correlation coefficient.
As shown in Fig.˜2 (Left), we observe a likelihood displacement phenomenon during optimization: the optimized model $\pi_{\theta}^{*}\text{ satisfies }P_{\pi_{\theta}^{*}}(y_{c}^{\prime}|x,c)<P_{\pi_{\theta}}(y_{c}^{\prime}|x,c)$ and $P_{\pi_{\theta}^{*}}(y_{p}|x,c)<P_{\pi_{\theta}}(y_{p}|x,c)$, i.e. probability mass is driven away from both the composite response $y_{c}^{\prime}$ and the parametric-based response $y_{p}$.
To understand where probability mass goes and ensure analytical tractability, we analyze the instantaneous update to the next-token distribution under gradient flow. Building on Theorem 5 of LD , let $W_{z}(t)$ denote the unembedding vector of token $z$ at training time $t$. The instantaneous change in the log-probability of token $z$, conditioned on input context $(x,c)$, is given by:
$\frac{d}{dt}\ln\pi_{\theta(t)}(z|x,c)\propto\langle W_{z}(t),W_{\text{token}(y_{c}^{\prime})}(t)-W_{\text{token}(y_{p})}(t)\rangle,$ (3)
Here, $W_{\text{token}(y_{c}^{\prime})}(t)$ is the unembedding vector of the token that the model is likely to generate given the context, while $W_{\text{token}(y_{p})}(t)$ corresponds to the token likely generated from the parametric-knowledge. In other words, the larger the inner product $\langle W_{z}(t),W_{\text{token}(y_{c}^{\prime})}(t)-W_{\text{token}(y_{p})}(t)\rangle$, the more positive the change in $\pi_{\theta(t)}(z|x,c)$.
Let $V(t)=W_{\text{token}(y_{c}^{\prime})}(t)-W_{\text{token}(y_{p})}(t)$ denote the direction vector. We analyze how different types of output tokens $z$ probability varied by examining the inner product $\langle W_{z}(t),V(t)\rangle$.
  • Faithful Token $z_{c}$ (derived from context $c$): With the LLM’s inherent ability to follow external context RAG ; zhou2023context ; gao2023retrieval , when generating the preferred response $y_{c}^{\prime}$ conditioned on $c$, the model is highly likely to produce tokens consistent with the context. Thus, $W_{\text{token}(y_{c}^{\prime})}$ is expected to be well aligned with $W_{z_{c}}$. In contrast, since $y_{p}$ reflects ungrounded, parametric-only generation, $W_{z_{c}}$ is likely misaligned with $W_{\text{token}(y_{p})}$. Therefore, the inner product $\langle W_{z_{c}}(t),V(t)\rangle$ is expected to be large.
  • Parametric Token $z_{p}$ (derived from internal knowledge, potentially hallucinated): The token $z_{p}$ is likely aligned with $W_{\text{token}(y_{p})}$, reflecting the model’s internal parametric memory. However, its alignment with $W_{\text{token}(y_{c}^{\prime})}$ is expected to be weak or negative. Consequently, $\langle W_{z_{p}}(t),V(t)\rangle$ is expected to be small.
  • Irrelevant Token $z_{\text{other}}$ (unrelated to context $c$ or parametric response $y_{p}$): $W_{z_{\text{other}}}$ is unlikely to exhibit strong alignment with either the context-dependent $W_{\text{token}(y_{c}^{\prime})}$ or the internal knowledge-based $W_{\text{token}(y_{p})}$. As a result, $\langle W_{z_{\text{other}}}(t),V(t)\rangle$ is expected to be small.
Let $\Delta P(z)$ denote the increase in probability for token $z$ due to likelihood displacement, proportional to $\frac{d}{dt}\ln\pi_{\theta(t)}(z|x,c)$. Based on the analysis of the alignment above, we have:
$\langle W_{z_{c}}(t),V(t)\rangle\gg\langle W_{z_{p}}(t),V(t)\rangle\quad\text{and}\quad\langle W_{z_{c}}(t),V(t)\rangle\gg\langle W_{z_{\text{other}}}(t),V(t)\rangle.$
Therefore, in the RAG setting, the likelihood displacement mechanism preferentially transfers probability mass towards tokens $z_{c}$ that are consistent with the external context $c$. This constitutes a benign likelihood displacement, actively promoting faithfulness by reinforcing context-aligned generation while suppressing tokens derived from parametric knowledge or irrelevant content.
Figure 3: Case study from the MemoTrap dataset illustrating benign likelihood displacement. The probability mass shifts from the parametric knowledge based token $z_{p}$ to the external knowledge based token $z_{c}$ after SSFO optimization.
3.4 Empirical Validation of Benign Likelihood Displacement
Setting. Our experiments utilize the MemoTrap dataset memotrap , designed to evaluate whether language models exhibit memorization traps. MemoTrap consists of instructions prompting the model to complete well-known proverbs with endings that deviate from the common completion. For instance, given the prompt "Write a quote that ends in the word ’right’: If you want a thing done right, do it __", the instructed target completion is "right". In this context, the token "right" represents external knowledge token $z_{c}$, while the commonly memorized completion token "yourself" is considered to be based on parametric knowledge token $z_{p}$.
We observe a benign form of likelihood displacement. As shown in Fig.˜2(Right), the probability of the faithful token $z_{c}$ increases after SSFO training. Furthermore, this rise is mirrored by a complementary fall in the probability of the parametric token $z_{p}$, producing a pronounced negative Pearson correlation (r=0.58r=-0.58). Probabilities for all remaining vocabulary tokens remain essentially unchanged and show no discernible correlation with $z_{c}$. A case study illustrating this displacement is presented in Fig.˜3.
3.5 Encouraging Benign Probability Displacement with SSFO-$\lambda$
As established in the previous analysis, the SSFO framework naturally induces a benign form of likelihood displacement, wherein probability mass shifts away from responses relying on parametric knowledge ($y_{p}$) toward those grounded in external context ($y_{c}$). This mechanism enhances model faithfulness by amplifying context-aligned generation. To further promote this desirable effect, we introduce SSFO-$\lambda$, a simple yet effective variant that explicitly encourages this displacement.
In contrast to prior work pal2024smaug ; dposhift ; AlphaPO ; Cal-dpo that focuses on mitigating likelihood displacement where $\pi_{\theta}(y_{w}\mid x)$ decreases, and even finds that encouraging the likelihood displacement may cause the model to exhibit crash behavior, producing unreadable responses dposhift . However, in Section˜3.3, we show that in the context-based question answering setting, a benign form of probability displacement can be encouraged to enhance the models’ faithfulness to the provided context through self-supervised preference optimization.
As proofed in dposhift , we introduce a parameter $\lambda>1$ to encourage the likelihood displacement during optimization:
$\mathcal{L}_{\text{SSFO}-\lambda}\left(\pi_{\theta};\pi_{\mathrm{ref}}\right)=-\mathbb{E}_{\left(x,c,y_{c}^{\prime},y_{p}\right)\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{c}^{\prime}\mid x,c\right)}{\pi_{\mathrm{ref}}\left(y_{c}^{\prime}\mid x,c\right)}-\lambda\cdot\beta\log\frac{\pi_{\theta}\left(y_{p}\mid x,c\right)}{\pi_{\mathrm{ref}}\left(y_{p}\mid x,c\right)}\right)\right].$ (4)
Figure 4: Correlation plot illustrating Span Exact Match scores for NQ-Swap, NQ-Open, and MemoTrap (scaled to the left y-axis) and ROUGE-L F1 scores for ELI5 (scaled to the right y-axis). Grey lines depict the regression trends. $r$ denotes Pearson correlation.
Empirical Validation of $\lambda$’s Effect. In Fig.˜4, we investigate the impact of varying $\lambda$ from 1.0 to 1.5 across multiple context-based question-answering benchmarks: NQ-Swap, NQ-Open, MemoTrap, and ELI5. As $\lambda$ increases we observe a consistent improvement in performance across all evaluated tasks. Pearson correlation coefficients $r$ reveal a positive relationship between $\lambda$ and performance on all datasets. For instance, span EM on NQ-Swap rises by about 1.6 points (from 77.2% to 78.8%), MemoTrap gains 1.7 points (from 74.8% to 76.5%). These results confirm that strategically amplifying the weight on the ungrounded (parametric) response via $\lambda>1$ indeed encourages a benign likelihood displacement and yields more faithful outputs.
Gradient Analysis. To further understand how this modification encourages the desired displacement, we analyze the gradient of this loss function. Following the derivation in dposhift , the gradient with respect to parameters $\theta$ is:
$\nabla_{\theta}\mathcal{L}_{\text{SSFO}-\lambda}=-\mathbb{E}\Big{[}c^{\prime}_{1}\big{(}\underbrace{\nabla_{\theta}\log\pi_{\theta}(y_{c}^{\prime}|x,c)}_{\text{increase likelihood of }y_{c}^{\prime}}-\underbrace{\lambda\nabla_{\theta}\log\pi_{\theta}(y_{p}|x,c)}_{\text{decrease likelihood of }y_{p}}\big{)}\Big{]},$ (5)
where $c^{\prime}_{1}$ is a positive coefficient. We present a detailed derivation of Eq.˜5 in Section˜C.1. Compared to the standard DPO update, Eq.˜4 applies a stronger negative weight ($\lambda$ where $\lambda>1$) to the gradient component associated with the parametric response $y_{p}$. Therefore, this parameter leads to a more pronounced suppression of the likelihood of the parametric response y during optimization.
4 Experiments
Datasets: To comprehensively evaluate faithfulness, we assess model performance across several dimensions.
Metrics: For short-form QA datasets (NQ-Open, NQ-Swap, MemoTrap, SQuAD), we adopt a standard zero-shot setting simulating a RAG scenario where the model answers queries based on the provided context without in-context learning’s demonstration. Performance is measured using the span Extraction Matching (span EM) score (a prediction is deemed correct if any segment of the generated output precisely matches one of the reference answers). For the long-form generation dataset ELI5, we report ROUGE scores rouge comparing the generated response against the ground-truth answer to assess content quality. For instruction following (FollowBench), we report Consistent Satisfaction Levels (CSL) followbench , which measures how many consecutive levels of instruction hardness a model can satisfy.
Models and Baselines: To ensure the generality of our approach, we conduct experiments using three families of open-source large language models LLaMA 3 touvron2023llama , Qwen 2.5 qwen2 and, Mistral as base models. We compare the proposed method against the strong publicly available methods focused on improving faithfulness: DECORE gema2024decore , ChatQA liu2025chatqa , Trust-Align trust_align , and Context-DPO context-dpo .
To construct the preference dataset, we randomly sample 900 query–context pairs from the MS MARCO datasetMS_MARCO , following the procedure outlined in Section˜3.2. Full implementation details are provided in Section˜D.4.
4.1 Faithfulness Evaluation Results
[…]
Instruct Model Method Implement
Human/Superior AI
Supervision
Robustness Response Quality
NQ-Swap Memo-Trap NQ-Open SQuAD Eli5
Span EM $\uparrow$ Span EM $\uparrow$ Span EM $\uparrow$ Span EM $\uparrow$ R-1 F1 $\uparrow$ R-2 F1 $\uparrow$ R-L F1 $\uparrow$
Llama-3-8B Base model \ \ 73.54% 73.60% 80.15% 88.20% 26.88% 9.12% 23.88%
Decoding-Stratagy DECORE 80.53% 74.40% 82.03% 84.90% 28.45% 10.96% 25.65%
Post-training ChatQA 67.70% 30.60% 76.80% 88.50% 28.61% 12.67% 26.26%
Trust-Align 75.56% 70.95% 77.38% 50.90% 11.36% 10.55% 20.15%
Context-DPO 82.76% 72.90% 82.86% 89.90% 28.15% 10.43% 24.98%
SSFO 81.23% 76.28% 84.40% 89.00% 30.40% 11.80%   27.01%
SSFO-$\lambda$  ✗ 82.81% 78.38% 85.69% 90.90% 31.35% 13.37%   28.37%
Qwen2.5-7B Base model \ \ 79.35% 54.19% 82.29% 90.30% 23.08% 6.06% 20.55%
Decoding-Stratagy DECORE 81.93% 54.56% 83.76% 82.80% 24.31% 7.01% 21.46%
Post-training Trust-Align 79.69% 53.71% 77.93% 80.30% 23.15% 6.45% 20.24%
Context-DPO 82.13% 55.34% 83.13% 91.80% 23.68% 6.53% 21.13%
SSFO 84.18% 57.66% 83.88% 92.00% 24.63% 7.17%   21.83%
SSFO-$\lambda$ 84.88% 60.77% 84.48% 93.30% 28.53% 10.40%   25.66%
Mistral-7B Base model \ \ 67.76% 34.34% 79.13% 84.80% 23.20% 5.74% 20.38%
Decoding-Stratagy DECORE 78.17% 30.68% 86.52% 85.30% 23.90% 6.48% 21.19%
Post-training Context-DPO 79.62% 33.20% 80.68% 86.50% 24.97% 7.02% 22.00%
SSFO 86.66% 37.22% 87.53% 89.00% 32.67% 13.47%   28.35%
SSFO-$\lambda$ 85.48% 46.91% 90.32% 88.50% 39.22% 21.36%   37.24%
SSFO and SSFO-$\lambda$ achieve state-of-the-art faithfulness across multiple datasets and models, as shown in Table˜1:
  • Both variants of Self-Supervised Direct Preference Optimization (SSFO and SSFO-$\lambda$) substantially improve contextual faithfulness over the base models. For example, SSFO shows an improvement of 5-20% regarding Robustness over parametric knowledge, highlighting the effectiveness of self-supervised faithfulness optimization regarding response quality. We report further evaluation results across models with varying parameter sizes in Appendix˜E.
  • SSFO-$\lambda$ attains state-of-the-art performance compared to both post-training-based baselines (ChatQA, Trust-Align, Context-DPO) and decoding strategy-based baseline (DECORE). These results confirm that explicitly encouraging benign likelihood displacement not only enhances faithfulness but also improves overall response quality.
[…]
Crose-language Response Quality Instruction Following
XQuAD(ES) DuReader(CN) FollowBench
Instruct Model Method Implement
Training Data
Required
Span EM $\uparrow$ Span EM $\uparrow$ CSL $\uparrow$
Base model \ \ 78.60% 78.80% 2.54
Decoding-Stratagy DECORE \ 83.87% 79.89% 2.46
ChatQA $\sim$30k 77.98% 72.05% 1.04
Trust-Align $\sim$15k 20.17% 8.12% 0.12
Context-DPO $\sim$5k 83.03% 84.40% 2.46
SSFO $\sim$800 83.10% 84.90%   2.70
Llama-3-8B Post-training SSFO-$\lambda$ $\sim$800 84.12% 83.56%   2.50
Base model \ \ 78.90% 81.50% 2.68
Decoding-Stratagy DECORE \ 80.08% 76.57% 2.56
Trust-Align $\sim$15k 75.21% 73.61% 0.58
Context-DPO $\sim$5k 79.92% 82.78% 2.64
SSFO $\sim$800 79.83% 83.27%   2.70
Qwen2.5-7B Post-training SSFO-$\lambda$ $\sim$800 81.76% 87.72%   2.62
Table 3: Faithfulness improvement case study: comparing the base model, Context-DPO and SSFO on equivalent English–Spanish (ES) query–answer pairs. This example from XQuAD artetxe2020cross demonstrates that SSFO improves faithfulness over the base model and preserves faithfulness cross-lingually. Blue texts highlight context-grounded information; Red texts mark unfaithful hallucinations.
Query (EN) What is Cultural Imperialism often referred to as?
(ES) ¿A qué se refiere a menudo el imperialismo cultural?
Context (EN)Cultural imperialism is when a country’s influence is felt in social and cultural circles, i.e. its soft power, such that it changes the moral, cultural and societal worldview of another…
(ES)El imperialismo cultural es cuando la influencia de un país se siente en los círculos sociales y culturales, es decir, su poder blando, de tal manera que cambia la cosmovisión moral, cultural y social de otro…
Base Model (EN) Cultural Imperialism is often referred to as cultural hegemony or cultural dominance. Hallucination
Context-DPO (EN) Cultural Imperialism is often referred to as the spread of a dominant culture’s influence. This is also known as the use of soft power.                               Partially Context-Faithful
(ES) El imperialismo cultural suele aludir a la influencia de la cultura de un país sobre la esfera sociocultural de otro, transformando su cosmovisión.                               Hallucination
SSFO (EN) A country’s influence is felt in social and cultural circles, i.e. its soft power, such that it changes the moral, cultural, and societal worldview of another… Context-Faithful
(ES) La influencia de un país se siente en los círculos sociales y culturales, es decir, su poder blando, de tal manera que cambia la cosmovisión moral, cultural y social de otro… Context-Faithful
4.2 Generalization Across Tasks and Languages
SSFO can enhance multi-language faithfulness. We evaluate the generalization ability of SSFO in Table˜2, and results show it can achieve cross-lingual faithfulness using only an English-based training set. For instance, on Llama-3-8B, SSFO increases Span EM scores by 6.10% on DuReader (Chinese) and 5.52% on XQuAD (Spanish) compared to the base model. In contrast, heavily supervised methods like ChatQA and Trust-Align exhibit decreased performance on these non-English QA datasets. This shows that by training the model to prioritize context knowledge over parametric knowledge, it learns a principle of contextual adherence that can transfer across languages.
SSFO minimally impacts instruction following capability. Requiring only a few hundred self-supervised data examples, SSFO largely preserves, and even slightly enhances, the model’s instruction following capabilities.The CSL scores on FollowBench indicate that models fine-tuned with SSFO retain comparable general instruction-following ability to the original base instruct models. We provide several cases that SSFO retains strong general instruction following ability under context-based scenarios in Table˜4. In contrast, other post-training approaches such as Trust-Align improve faithfulness at the cost of degrading general generative abilities (e.g., CSL dropping to 0.12 for LLaMA-3-8B).
Table 4: Case study from FollowBench followbench : showing that SSFO retains strong instruction-following capabilities under a complex, context-based NLP task.
Instruction Analyze the provided text to pinpoint a category and the sentiment (positive, neutral, or negative) it emanates. Additionally, perform named entity recognition to emphasize notable entities and also identify the core topic discussed. Select the category from this array: company, educational institution, artist, athlete, office holder, means of transportation, building, natural place, village, animal, plant, album, film, or written work.
Context Michael DenDekker - Michael G. DenDekker (born July 11, 1961) is an assemblyman for the state of New York’s 34th district which includes the neighborhoods of Woodside, Jackson Heights, and East Elmhurst, all in the borough/county of Queens.
Trust Align I apologize, but I couldn’t find an answer to your question in the search results.
SSFO Category: Office Holder; Sentiment: Neutral ; Notable entities: Michael DenDekker, State of New York, 34th district, Woodside… Core topic discussed: Michael DenDekker’s role as an assemblyman for the state of New York’s 34th district.
Rationale: Note that the text does not contain any information about the office holder’s achievements, opinions, or personal life, which is why the sentiment is neutral… General Ability Retained
4.3 Data Efficiency Analysis
Figure 5: Data efficiency study: SSFO requires about 60% of data (400–500 examples) to achieve 85% of the total possible performance gain over the base model.
To measure how many self-supervised preference examples SSFO actually needs, we subsample the preference dataset in 10% increments. As shown in Fig.˜5, we evaluate the average performance gain (an average improvement over the base instructed model). Both SSFO models cross the 85% performance threshold by approximately 50–60 % of the data ( 400–500 examples). We hypothesize that this efficiency stems from SSFO’s ability to leverage data generated by the model itself. This self-generated data inherently matches the model’s response style, thereby eliminating stylistic discrepancies and allowing the model to concentrate on optimizing the faithfulness gap. We compare the training examples from SSFO with other post-training methods in Section˜D.3.
5 Limitation
This paper introduces SSFO, a novel and efficient self-supervised method that significantly enhances faithfulness in RAG systems without inference overhead. While the current work demonstrates strong performance, one potential avenue for future investigation could be to extend its application and further assess its efficacy in more complex RAG scenarios, such as those involving extremely long contexts or the synthesis of information from multiple retrieved documents. Investigating the scalability and adaptability of SSFO to even more intricate and diverse knowledge domains represents a direction for continued advancement.
6 Conclusion
This work addressed the critical challenge of faithfulness hallucination in Retrieval-Augmented Generation (RAG) systems, where existing methods often introduce significant computational overhead or rely on costly external supervision. We introduced Self-supervised Faithfulness Optimization (SSFO), a novel and efficient self-supervised alignment method. SSFO uniquely leverages the LLM’s own outputs—contrasting responses generated with external contexts against those relying solely on parametric knowledge—to create preference pairs for alignment. Our experiments across diverse benchmarks show that SSFO significantly enhances model faithfulness and robustness against parametric knowledge, achieving state-of-the-art performance compared to existing methods. Furthermore, SSFO exhibits strong generalization capabilities, improving faithfulness even in cross-lingual settings using only English training data, while crucially preserving the model’s general instruction-following abilities.
Appendix A Related Work
A.1 Faithfulness Hallucination of Large Language Models
Hallucination in LLMs can be generally categorized into two types: factuality hallucination, where generated content deviates from established world knowledge (e.g., claiming “Mars has oceans”), and faithfulness hallucination, where the generated response is inconsistent with the provided context(e.g., misrepresenting a source document’s information). [9]
Current methods to address faithfulness hallucination primarily fall into two categories:
  • Post-training-based methods rely on supervised fine tuning [4, 41] or preference alignment [13]. [12] propose a two-stage instruction tuning method and create a dataset (including human annotation) that aiming at enhancing LLM’s capability of integrating external context. For alignment based methods, one key factor lies in create pairwise preference dataset: [5] use GPT-4 [3] to generated well cross referenced response as positive answer and use Llama2 [4] to generate negative response results in an alignment dataset of 15K samples . In [6], the author prompts GPT-4 to correct hallucinations in the response which use as positive response and the original response as negative one. Context-DPO creates preference data by perturbing a knowledge graph and employ GPT-4 to generated counterfactual context [11]. While these methods can yield more customized responses, they often demand costly supervision from humans or advanced AI models and can lead to extensive post-training processes that may cause catastrophic forgetting [18, 42], thereby undermining the model’s generalization capabilities
  • decoding strategy-based methods: In [15] the author present context-aware decoding (CAD), which follows a contrastive output distribution that amplifies the difference between the output probabilities when a model is used with and without context. DECORE [14] extend this framework to masking retrieval heads to induce faithfulness hallucinations, followed by a dynamic entropy-controlled contrastive decoding to penalize uncertain outputs. While these methods are training-free and adaptable[cite: 234], they often significantly increase the inference burden, typically by requiring parallel processing
To overcome these challenges, this paper introduces Self-Supervised Faithfulness Optimization (SSFO), a novel self-supervised alignment method that enhances faithfulness without introducing external supervision or additional inference burden.
A.2 Direct Preference Optimization and Likelihood Displacement
RLHF [43, 44] requires fitting a reward model to a dataset of human (or AI) preferences, and then training the language model to maximize the reward, which is computationally expensive and can suffer from instabilities. This has led to the rise of direct preference optimization (DPO) [13] . DPO implicitly optimizes the same objective as RLHF algorithms but is simple to implement and straightforward to train.
Likelihood displacement [17] refers to the counterintuitive phenomenon where during direct preference alignment, while the gap between preferred responses and unpreferred response increases, they both decrease. Such a phenomenon is unwanted since the preferred response other derivative from human annotator or strong AI model. To alleviate this problem, DPOP [23] design a modified DPO loss function to penalizes reducing the probability of the positive completion; AlphaPO [26] introduce a parameter to adjust the shape of the reward function beyond standard log rewards, providing fine control over the likelihood displacement; DPO-Shift [25] adds a real-valued function to controllably shift the distribution of the preferred probability.
Existing approaches typically assume the preferred response is a "golden" label and aim to alleviate likelihood displacement. In contrast, SSFO optimizes the model using self-supervised preference data, which can be considered "silver" labels, yet still provide a clear supervisory signal towards faithfulness. This work demonstrates that, in the RAG setting, likelihood displacement can be a benign phenomenon and can even be encouraged to benefit the faithfulness alignment process.
Appendix B Broader impacts
Our work on Self-Supervised Faithfulness Optimization (SSFO) presents significant positive impacts for the advancement of reliable and trustworthy AI systems. By enhancing the faithfulness of Retrieval-Augmented Generation (RAG) models to provided context, SSFO alleviate the critical issue of faithfulness hallucination. This improvement leads to more accurate, verifiable, and dependable outputs from Large Language Models, which is crucial for applications where information integrity is paramount, such as in educational tools, scientific research, and systems providing critical information to the public. Reducing the propensity of LLMs to generate content that deviates from factual sources fosters greater user trust and promotes the responsible deployment of AI technologies in diverse real-world scenarios.
Furthermore, the self-supervised nature of SSFO offers considerable practical benefits that can accelerate the adoption of more faithful AI. By eliminating the need for costly human annotation or extensive, resource-intensive post-training procedures, our method makes the development of highly faithful models more accessible and economically viable for a broader range of researchers and developers. The negligible inference burden and the demonstrated strong generalization capabilities, including improved cross-lingual faithfulness and preservation of general instruction-following abilities, mean that the benefits of SSFO can be widely applied across different languages and tasks. This facilitates the development of more robust and equitable AI systems globally, contributing to a more informed and reliably assisted digital environment.
Appendix C Mathematical Derivations
C.1 Gradient Derivation for $\mathcal{L}_{\text{SSFO}-\lambda}$
The loss function for $\lambda$-DPO is given by:
$\mathcal{L}_{\text{SSFO}-\lambda}(\pi_{\theta},\pi_{ref})=-\mathbb{E}_{(x,c,y_{c}^{\prime},y_{p})\sim\mathcal{D}_{pref}}\left[\log\sigma(u)\right]$ (6)
where
$u:=\beta\log\frac{\pi_{\theta}(y_{c}^{\prime}|x,c)}{\pi_{ref}(y_{c}^{\prime}|x,c)}-\lambda\cdot\beta\log\frac{\pi_{\theta}(y_{p}|x,c)}{\pi_{ref}(y_{p}|x,c)}$ (7)
The gradient with respect to $\theta$ is:
$\nabla_{\theta}\mathcal{L}_{\text{SSFO}-\lambda}=-\mathbb{E}\left[\frac{\sigma^{\prime}(u)}{\sigma(u)}\nabla_{\theta}u\right]$ (8)
Using the properties of the sigmoid function $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))$ and substituting u-u, the gradient simplifies to:
$\nabla_{\theta}\mathcal{L}_{\text{SSFO}-\lambda}$ $=-\mathbb{E}\left[\sigma\left(\lambda\cdot\beta\log\frac{\pi_{\theta}(y_{p}|x,c)}{\pi_{ref}(y_{p}|x,c)}-\beta\log\frac{\pi_{\theta}(y_{c}^{\prime}|x,c)}{\pi_{ref}(y_{c}^{\prime}|x,c)}\right)\right$.
$\quad\left.\times\left(\beta\nabla_{\theta}\log\pi_{\theta}(y_{c}^{\prime}|x,c)-\lambda\cdot\beta\nabla_{\theta}\log\pi_{\theta}(y_{p}|x,c)\right)\right]$ (9)
Let $c^{\prime}_{1}$ be defined as:
$c^{\prime}_{1}:=\beta\sigma\left(\lambda\cdot\beta\log\frac{\pi_{\theta}(y_{p}|x,c)}{\pi_{ref}(y_{p}|x,c)}-\beta\log\frac{\pi_{\theta}(y_{c}^{\prime}|x,c)}{\pi_{ref}(y_{c}^{\prime}|x,c)}\right)$ (10)
Then the final gradient form is:
$\nabla_{\theta}\mathcal{L}_{\text{SSFO}-\lambda}=-\mathbb{E}\left[c^{\prime}_{1}\left(\nabla_{\theta}\log\pi_{\theta}(y_{c}^{\prime}|x,c)-\lambda\nabla_{\theta}\log\pi_{\theta}(y_{p}|x,c)\right)\right]$ (11)
This matches the target formula.
Appendix D Additional Details
D.1 Datasets
We utilize a variety of datasets to comprehensively evaluate the proposed SSFO method across different aspects of faithfulness, response quality, and generalization to ensure consistently strong performance in a retrieval-augmented generation (RAG) setting.
  • MemoTrap [29] is designed to reveal “memorisation traps” by pitting a well-known proverb against a context-correct but counter-habitual ending. Example (prompt): “Write a quote that ends in the word ‘right’: If you want a thing done right, do it    ” – the context expects the completion right, not the cached continuation yourself
  • Open Book NQ-Open (NQ-Open) [33] is an open-domain QA benchmark provide with supporting passages. Example: Passage: Vatican City…. is the smallest country in Europe by both area and population; Question: Which country has the smallest population in europe?” → Vatican City
  • NQ-Swap [32] extends Natural-Questions with entity swaps to create conflicts between retrieved context and parametric memory. Example: Context states “Ferraro is known for her portrayal of Grace Bowman in The Secret Life of the American Teenager”; the query asks “Who plays Grace in …?” – the correct answer is Ferraro, although parametric knowledge often yields Molly Ringwald
  • SQuAD v1.1 [34] provides short passages with span-based questions. Example: Passage: “Google was founded in 1998 by Larry Page and Sergey Brin”; “Question: Who founded Google?” → Larry Page; Sergey Brin
  • ELI5 [35] contains long-form, lay-audience explanations. Example: “Why is the sky blue?” expects a multi-sentence answer discussing Rayleigh scattering.
  • DuReader (Chinese) [36] evaluates cross-lingual comprehension with Web passages. Example (in English for illustration): “Who wrote Dream of the Red Chamber?” → Cao Xueqin.
  • XQuAD (Spanish split) [37] probes zero-shot transfer to non-English languages. Example (in English for illustration): “Who was the first person to transmit radio waves across the Atlantic?” → Guglielmo Marconi.
  • FollowBench [38] measures fine-grained instruction following. Example: To enhance your time management skills, can you devise a method incorporating a mind map and featuring a touch of alliteration in the suggestion, ensuring each sentence contains no more than 15 words?
D.2 Baselines
Instruct Model [4, 40]: A vanilla instruction-tuned LLM queried with a standard retrieval-augmented generation (RAG) prompt. No additional alignment is applied.
Context-DPO [11] : Improves context faithfulness by applying Direct Preference Optimization on the ConFiQA benchmark, which injects knowledge conflicts to mimic real RAG scenarios.
Trust-Align [15]: Builds GPT-4 “gold” answers cross-referenced to the retrieved context as positive samples and Llama outputs as negative samples, then performs DPO to steer the model toward faithful responses.
DECORE [14]: A training-free decoding strategy that reduces hallucinations by contrasting outputs of the base LLM and a masked variant (retrieval heads suppressed) guided by conditional entropy.
ChatQA [12]: Enhances RAG and conversational QA via a two-stage instruction tuning method and a dense retriever optimized for dialogue, reducing deployment costs while matching query rewriting models.
D.3 Details on Self-supervised Preference data construction
Starting from the MS MARCO [27] passage-ranking corpus, we first draw 900 $(\textit{query},\textit{context})$ pairs at random, ensuring topic diversity and open-domain coverage. We use prompts in Table˜5 to construct preferred and unpreferred response from the base-instruct model.
This yields a self-supervised dataset in which both strings share the model’s native style. After discarding empty answers (unpreferred response is I do not know) we keep $966/643/842$ triplets for Llama-3-8B, Qwen-2.5-7B and Mistral-7B, respectively. In practice, we train the base model $\sim 500$ pairs already deliver $\geq\!85\%$ of the final gain.
Table 5: Prompts for self-supervised preference data generation.
Prompt for preferred response
Based on the following context:
Context: {Context}
Question: {Question}
If you are not sure of the answer, please reply ’I don’t know’.
Prompt for unpreferred response
Question: {Question}
If you are not sure of the answer, please reply ’I don’t know’.
We compare generated preference data in Table˜6, Trust-Align induces a large style gap—the preferred response is verbose and citation-driven, while the unpreferred one is concise and citation-light; Context-DPO performs closed-form QA with simple entity swaps (e.g., “microfiber cloth” vs. “paper towel”); SSFO (Open-Domain QA) preserves the model’s native style for both preferred and unpreferred answers, enabling optimization to focus purely on faithfulness improvements without introducing a confounding style gap.
Table 6: Comparison of response generated by different post-training based methods.
Strategy Preferred Response Unpreferred Response
Trust-Align To safely clean your computer screen, first power off the device, as this makes dust and smudges easier to see and reduces the risk of damage [2][3]. Use only a clean, dry microfiber cloth as microfiber is specifically designed to clean screens gently without scratching them [1][3]. First, turn off your laptop to prevent any damage to the screen. According to a guide by [1] How to Safely Clean Your Screen, this is essential for effective cleaning. …. References: [1] How to Safely Clean Your Screen
Context-DPO Lightly wipe the screen with a clean, dry microfiber cloth. Lightly wipe the screen with a clean, dry paper towel.
SSFO Based on the passages, to safely clean your computer screen:
1. Wipe gently with a microfiber cloth, avoiding excessive pressure.
2. Do not use paper towels, kitchen towels, or household rags.
To safely clean your computer screen, you can use a soft, dry cloth or a microfiber cloth to gently wipe it. Avoid liquids, chemicals, or abrasive materials that may damage the surface.
D.4 Implementation Details
Training and Inference settings.
Following Section˜3.1, we create a self-supervised preference set per model: 966 pairs for Llama-3-8B-Instruct, 643 for Qwen2.5-7B-Instruct, and 842 for Mistral-7B-Instruct-v0.3. We fine-tune each model with SSFO using a global batch size of 8 on 8×NVIDIA A6000 GPUs in BF16. Optimisation employs AdamW ($\beta_{1}{=}0.9,\;\beta_{2}{=}0.95$) with a peak learning rate of $1{\times}10^{-6}$, linearly warmed up from 0 over the first 10 steps and then held constant. For SSFO-$\lambda$, we set $\lambda=1.5$ for Llama-3 and Mistral, and $\lambda=1.3$ for Qwen2.5.
During both dataset generation and evaluation we decode with a temperature of 0.7.
Appendix E Additional Results
E.1 Effect of Model Scale on SSFO Performance
Table 7: Impact of SSFO on robustness and response quality across varying LLM scales (1.5B–72B parameters).
Instruct Model Method Robustness Response Quality
NQ-Swap Memo-Trap NQ-Open SQuAD Eli5
[…]
Qwen2.5 1.5B Base model 65.78% 25.44% 79.17% 88.80% 28.11% 10.26% 25.40%
SSFO 79.65% 41.12% 83.88% 92.90% 34.10% 14.86% 30.87%
Qwen2.5 3B Base model 76.38% 47.66% 76.95% 88.20% 23.57% 6.60% 21.11%
SSFO 82.11% 59.12% 81.32% 92.40% 25.74% 9.16% 23.31%
Qwen2.5 7B Base model 79.35% 54.19% 82.29% 90.30% 23.08% 6.06% 20.55%
SSFO 84.18% 57.66% 83.88% 92.00% 24.63% 7.17% 21.83%
Qwen2.5 14B Base model 82.15% 64.08% 82.49% 90.00% 22.48% 5.62% 20.03%
SSFO 85.04% 66.52% 84.14% 92.80% 25.42% 8.19% 22.97%
Qwen2.5 72B Base model 79.84% 66.99% 83.50% 91.20% 21.85% 5.26% 19.45%
SSFO 82.51% 67.39% 84.48% 91.70% 22.46% 5.71% 19.96%
To assess the scalability of SSFO, we apply it to Qwen2.5 models with sizes from 1.5 B to 72 B parameters and report relative gains over each base model on five benchmarks (Table˜7). Across all model sizes, SSFO consistently improves robustness—yielding +3 % to +15 % span EM gains—and enhances response quality—boosting closed-book QA and long-form generation metrics by up to +6 %. The largest relative improvements occur on smaller models (e.g., +13.9 % EM on Memo-Trap for Qwen2.5 1.5 B), while even the 72 B model sees steady gains (e.g., +2.7 % EM on NQ-Swap).
These results demonstrate that SSFO delivers a stable, scalable enhancement to faithfulness and answer quality across a wide spectrum of LLM sizes.
E.2 Instruct following
Table 8: Constraint Satisfaction Levels (CSL↑) on FollowBench across Llama-3-8B and Qwen2.5-7B models. Results are broken down by Content, Situation, Style, Format, and Mixed constraint categories, as well as their overall mean.
Instruct Model Method        Implement Supervision Instruction Following (FollowBench)
[…]
Llama-3-8B Base model \ \ 2.6 2.4 3.3 2.9 1.5 2.54
Decoding-Strategy DECORE 2.4 2.0 3.1 3.3 1.5 2.46
Post-training ChatQA 1.3 1.0 1.0 1.1 0.8 1.04
Trust-Align 0.1 0.1 0.0 0.0 0.4 0.12
Context-DPO 2.5 2.5 3.0 2.8 1.5 2.46
SSFO 2.8 2.7 3.2 3.2 1.6 2.70
SSFO-$\lambda$ 2.3 2.4 3.1 3.1 1.6 2.50
Qwen2.5-7B Base model \ \ 2.6 3.5 2.9 2.8 1.6 2.68
Decoding-Strategy DECORE 2.4 3.2 2.7 2.8 1.7 2.56
Post-training Trust-Align 0.5 1.5 0.3 0.1 0.5 0.58
Context-DPO 2.6 3.3 3.0 2.9 1.4 2.64
SSFO 2.6 3.1 3.2 2.9 1.7 2.70
SSFO-$\lambda$ 2.6 3.0 3.2 2.7 1.6 2.62
In Table˜8, we present detailed sub-metrics for several faithfulness-enhancement methods to evaluate their impact on LLMs’ overall instruction-following capabilities. Among these approaches, SSFO not only best preserves general instruction adherence but also delivers the gains on Content-category tasks—a result we attribute to SSFO’s ability to steer the model to attend more closely to the source text. SSFO emerges as a practical technique for boosting faithfulness while maintaining the model’s effectiveness on general-purpose tasks. For example, in a RAG-powered law assistant, in a legal research setting, SSFO preserves verbatim case-law citations yet allows users to reorganize, highlight, and summarize key holdings according to specified style and formatting instructions.
E.3 Case Study
We provide more qualitative illustration of SSFO’s benefits. As shown in Table˜9 SSFO corrects parametric hallucinations made by the base model. In Table˜10, example from FollowBench highlights SSFO’s ability to retain strong instruction-following capabilities under complex, context-based scenarios. In Table˜11 a cross-language case study underscores SSFO’s capacity to maintain faithfulness even when queries and context are presented in a different language, such as Spanish.
Table 9: A case study showing how SSFO improves faithfulness on a NQ-swap dataset [32]. blue text indicates the correct, context-based answer; red text denote parametric-induced faithfulness hallucination.
Question Who plays grace in The Secret Life of the American Teenager?
Context Ferraro (born July 24, 1986) is a Canadian actress and singer. She is known for her portrayal of Grace Bowman in the television series The Secret Life of the American Teenager.
Base According to the text, Molly Ringwald plays the role of Grace Bowman in the television series The Secret Life of the American Teenager. Hallucination
SSFO Based on the text, Ferraro plays the role of Grace Bowman in the television series The Secret Life of the American Teenager. Context-Faithful
Table 10: Case study from FollowBench [38]: showing that SSFO retains strong instruction-following capabilities under a complex, context-based NLP task.
Instruction Below we show examples of soft drink. Following these examples, keep listing three more soft drink that satisfy following conditions: 1. not manufactured by coca cola company; First Three Examples:
Examples a) in chemistry, a phosphate is an anion, salt, functional group or ester derived from a phosphoric acid. b) coca cola, or coke, is a carbonated soft drink manufactured by the coca cola company.c) an egg cream is a cold beverage consisting of milk, carbonated water, and flavored syrup (typically chocolate or vanilla).
ChatQA D) Fanta E) Pepsi Cola F) Mecca cola Fanta is manufactured by coca cola company
SSFO Based on the given examples, here are three more soft drinks that satisfy the condition of not being manufactured by the Coca-Cola company: 1. Dr Pepper: a carbonated soft drink that is a blend of vanilla and cherry flavors. 2. Mountain Dew: a sweet and citrus-flavored carbonated soft drink that is owned by PepsiCo. 3. A&W Root Beer: a creamy and sweet carbonated soft drink that is flavored with vanilla and anise, owned by Dr Pepper Snapple Group (now Keurig Dr Pepper).                                         General Ability Retained
Table 11: Cross-language case study demonstrating that SSFO maintains faithfulness even when the query–answer pair is presented in Spanish (ES). Blue spans highlight correct, context-grounded information; Red spans mark parametric hallucinations.
Query (EN) Who first sent radio waves across the Atlantic?
(ES) ¿Quién fue el primero en enviar ondas de radio a través del Atlántico?
Context (EN)Guglielmo Marconi made his famous trans-Atlantic radio transmission for the first time in 1901… (193 words omitted)
(ES) …Cuando Guglielmo Marconi realizó por primera vez su famosa transmisión transatlántica de radio en 1901
Context-DPO (EN) According to the context, Guglielmo Marconi made the first-ever trans-Atlantic radio transmission in 1901.                                         Faithful to context
(ES) Según el contexto, no se puede determinar con certeza quién fue el primero en enviar ondas de radio a través del Atlántico, ya que la Corte Suprema de los Estados Unidos declaró que la decisión no tenía nada que ver con la reivindicación de Marconi… Hallucination
SSFO (EN) Based on the context, Guglielmo Marconi made the first-ever trans-Atlantic radio transmission in 1901.                                         Faithful to context
(ES) Según el contexto, Guglielmo Marconi fue el primero en enviar ondas de radio a través del Atlántico, en 1901.                                         Faithful to context
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: As shown in the abstract and the end of the introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discussed the limitations in Section˜5
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide the proof/derivation in the Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We state the experiment setting before the results and in appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer:[Yes]
  24. Justification: We provide the code in an anonymous link.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We state the experiment setting before the results and in appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provide statistical significance in appendix.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provided these information in implementation details.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We follow the Code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We state our work has the potential to promotes the responsible deployment of Retrieval Augmented Generation.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We use open-source datasets.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited corresponding works in the paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our code in the anonymous link has a detailed readme.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The method development in this research does not involve LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.